OpenStudy (anonymous):

graph each equation. 1.y=|x|-2 2.y=|x|+3 3.y=|x-3|+1 4.y=|x+1|-4

OpenStudy (jdoe0001):

y = A | Bx + C | + D C = horizontal shift, if C is negative, to the right, if C is positive, to the left D = vertical shift, if D is positive, up, if D is negative, down

OpenStudy (anonymous):

Equation 1: y=x-2 y intercept = -2 x intercept (y=0): 0=x-2 x=2 draw a line between the two points! 

OpenStudy (anonymous):